1. Suppose $26002600 is invested annually into an annuity that earns 55​% interest. If P dollars are invested annually at an interest rate of​ r, the value of the annuity after n years is given by the following equation. Upper W equals Upper P left bracket StartFraction left parenthesis 1 plus r right parenthesis Superscript n Baseline minus 1 Over r EndFraction right bracketW=P(1+r)n−1r

How much is the annuity worth after 5 years?

2. Suppose that ​$90,000 is invested at 66​% interest. Find the amount of money in the account after 4 years if the interest is compounded annually?
Answer #1

Answer:

Question 1)

Future value of Annuity FV=P*{(1+r)^n-1)/r}

P=Annuity payment =$26002600

r=Interest rate=55%

n=5 years

FV=26002600*{(1+55%)^5-1)/55%}

FV=$375,695,478.29

Question 2)

Future value of investment FV=P*(1+r)^n

P=Investment =$90000

r=Interest rate=66%

n=5 years

FV=90000*(1+66%)^5=$1,134,443.71